OpenStudy (anonymous):

I Need help from someone good at Math ASAP! Whoever does will get a good testimonial and MEDAL!!!!

OpenStudy (ash2326):

Wooho please post your question

OpenStudy (anonymous):

ok its a project

OpenStudy (anonymous):

Using the function A(t) = P(1 + r)t, create the function that represents your new car loan. The principle will be the price of the vehicle you selected, not how much you are putting down. Use your function to determine how much you will owe the bank in five years. Remember that the exponent t is measured in months. This is assuming you are not paying down your loan, so do not get scared! Show all of your work. Again, we are assuming that you are not paying down your loan, and the bank's interest keeps piling on. Being a smart financial planner, you want to figure out how many months it will be until you owe the bank an amount of $500,000.00. Solve for t and show all of your work. Lastly, you decide to keep track of your loan weekly instead of monthly. Solve for the adjusted interest rate. @ash2326

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

That should be easy A= amount of the money you owe the bank after t months ok choose the car and tell me its price? that will be P

OpenStudy (anonymous):

$114,395 @ash2326

OpenStudy (ash2326):

ok so we have \[A=P(1+r)*t\] P=114395 \[A=114395(1+r)*t\] now do you have the rate of interest?

OpenStudy (anonymous):

1.4% each month

OpenStudy (ash2326):

ok 1.4% which is 0.014 in decimal Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

okay, put these values in the equation and what will you put for t?

OpenStudy (anonymous):

what will i put?

OpenStudy (ash2326):

you know that period is 5 years, so t?

OpenStudy (anonymous):

5 right?

OpenStudy (ash2326):

read the second paragraph again

OpenStudy (ash2326):

woohoo give me some time, I will finish dinner and come back. meanwhile you try this part

OpenStudy (anonymous):

500,000

OpenStudy (ash2326):

t would be in months

OpenStudy (anonymous):

so yeah 5 right?

OpenStudy (ash2326):

come on? 5 is years

OpenStudy (anonymous):

oh wait 9 weeks right

OpenStudy (ash2326):

1 year = 12 months 5 years= ?

OpenStudy (anonymous):

it sasys to turn 5 months into week

OpenStudy (ash2326):

that's the last part of the question

OpenStudy (ash2326):

brb

OpenStudy (anonymous):

noooooo

OpenStudy (anonymous):

dont leave!!

OpenStudy (ash2326):

wait for sometime I'll be back,

OpenStudy (anonymous):

its not that i will be leaving for work and im going to be back after 6 hours can you jst tell me how to do it and then leave!?

OpenStudy (ash2326):

for the first part t=60 months ( 5years) Use this in the formula to find A

OpenStudy (anonymous):

ok 6959791.8

OpenStudy (anonymous):

is that right

OpenStudy (ash2326):

yes,

OpenStudy (anonymous):

ok now what do i do?

OpenStudy (ash2326):

now you have A and you need to find t A=500,000 use the same equation to find t?

OpenStudy (anonymous):

wasnt t 60 thats what i pllugged in the first equation

OpenStudy (ash2326):

that was for the first part, but now your loan amount has to increase to 500,000 and we need to find the time? obviously it'll be less than 5 years

OpenStudy (anonymous):

im leaving but pls leave the steps i have to turn this in today ill be back at 6

OpenStudy (ash2326):

ok, I'll see what I can do

OpenStudy (phi):

how far did you get ?

OpenStudy (anonymous):

just this far.

OpenStudy (phi):

ok, for this part: Using the function A(t) = P(1 + r)t, create the function that represents your new car loan. The principle will be the price of the vehicle you selected, not how much you are putting down. what is P ? what is r ?

OpenStudy (anonymous):

t is 50 months and p is the price of my car

OpenStudy (phi):

What number did you pick for P and what number for r ?

OpenStudy (anonymous):

p is 141,395 and r is 0.014

OpenStudy (phi):

The principle will be the price of the vehicle you selected That seems expensive for a car. Are you sure you want to use such a big number ?

OpenStudy (anonymous):

yeah its the price of the car im planning to get porsche

OpenStudy (anonymous):

should we change it?

OpenStudy (phi):

r is the interest rate. Normally we talk about interest rate for a year then, if we want to change the time t to months, we divide r by 12 (if we change t to weeks, we divide r by 52) what is the interest rate for the year ?

OpenStudy (anonymous):

im confused honestly

OpenStudy (phi):

Currently, interest rates are about 3% annual percentage rate or 0.03 they want you to find A(t) when t is in months. that means we make r= 0.03/12 or r= 0.0025

OpenStudy (phi):

now fill in the numbers for P and r in \[ A(t) = P(1 + r)^t, \]

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

whats t?

OpenStudy (phi):

for the moment, leave t as "t" we are making an equation as a function of t.

OpenStudy (anonymous):

ok hold up im going to solve it

OpenStudy (anonymous):

0.030075

OpenStudy (anonymous):

is that right?

OpenStudy (phi):

we are not looking for a number (yet) we want an equation that looks like \[ A(t) = P(1+r)^t \] but with P and r replaced with numbers (and leaving "t" as a letter)

OpenStudy (anonymous):

ok so you want me to give you the equation

OpenStudy (phi):

yes. I would use P=141,395 and r= 0.0025

OpenStudy (anonymous):

oh ok a(t)=141,395(1+0.0025)^t

OpenStudy (anonymous):

@phi

OpenStudy (phi):

looks good. notice you can add 1+0.0025 and get 1.0025 so your equation is \[ A(t) = 141,395\cdot 1.0025^t\] where t is measured in months

OpenStudy (anonymous):

50

OpenStudy (phi):

Next ***Use your function to determine how much you will owe the bank in five years. *** how many months are in 5 years ?

OpenStudy (anonymous):

60 months

OpenStudy (phi):

so they want you to find A(60)

OpenStudy (anonymous):

you mean +1.0025 you put multiply

OpenStudy (phi):

you evaluate \[ A(60)= 141,395\cdot 1.0025^{60} \] remember your order of operations. First do the exponentiation. then multiply by 141395.00

OpenStudy (anonymous):

ok, 45.624112452224 @phi

OpenStudy (phi):

what did you get for 1.0025^60 ?

OpenStudy (anonymous):

1.161616781555

OpenStudy (anonymous):

@phi

OpenStudy (phi):

now what do you get after you multiply that times P?

OpenStudy (anonymous):

164246.8048

OpenStudy (phi):

ok, but this is money, so people round to the nearest cent. you get $164246.80 that is the answer to the first part.

OpenStudy (anonymous):

oh thanks! what the second part asking

OpenStudy (phi):

second part: you want to figure out how many months it will be until you owe the bank an amount of $500,000.00 in other words A(t)= 500,000 first step: write down the equation. Don't solve it, just write it down.

OpenStudy (anonymous):

0.03(1+0.0025)^500000?

OpenStudy (phi):

no. start with your equation that you figured out in the first part, \[ A(t) = 141,395\cdot 1.0025^t \] A(t) is how much you owe. A(t) is 500,000 replace A(t) with 500,000

OpenStudy (anonymous):

\[500,000=141,395*1.0025^t\]

OpenStudy (anonymous):

@phi

OpenStudy (phi):

now that you have the correct equation for this part, we solve for t. the first step is divide both sides by 141395

OpenStudy (anonymous):

3.536192935=1.0025^t

OpenStudy (phi):

now take the log (base 10) of both sides

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you write log( ) around each term. can you do that ?

OpenStudy (anonymous):

i got .5485359521

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you should always write the whole equation, not a random number.

OpenStudy (anonymous):

idk how to write that in the calculator thats the problem

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you don't use the calculator at every step. you should do this: \[ 500,000=141,395*1.0025^t \\ 3.536192935=1.0025^t \] take the log base 10 of both sides \[ \log_{10}(3.536192935) = \log_{10}(1.0025^t) \] on the right side, use log(a^b) = b*log(a) to rewrite it.

OpenStudy (phi):

can you simplify \[ log_{10}(1.0025^t) \] using \[ \log_{10}(a^t) = t \log_{10}(a) \]

OpenStudy (anonymous):

@SolomonZelman how would you type that in a calculator?